Sanjay Bhatia, (age 30 years) works with a pharmaceutical company and has not yet started to invest for his retirement. Sanjay is married to Preeti (age 28 years) and they have one son aged 2 years. Sanjay wants you to prepare a plan for him to retire latest by age 55. (You can make any assumptions to further build up your case.


1
Expert's answer
2020-12-13T17:19:31-0500

Solution to question number #150130

Between now(30 years) and age 55, Sanjay Bhatia has 25 years to invest for his retirement. Sanjay Bhatia can invest one time amount or a chosen amount for a specific period for 25 years.

Let us assume the target amount at year 55 is ₹20,000,000.

He can invest a one time amount which will grow to ₹20,000,000 for instance:

Amount = Future value/PVIF, n=25, r=6.5

PVIF = 1-(1+r)n-1/r=12.1979

Amount = ₹20,000,000/12.1979

Amount to invest now= ₹1,639,626.50 to get a return of 6.5% per annum.

Second option

A=FV/PVIFA

FV= 20,000,000 while PVIFA using ((1+r)-n-1/r)(1+r)=62.7154

so, A=₹20,000,000/62.7154

A=₹318,901

Invest an amount of ₹318,901 at the beginning of every year for 25 years to get a lump sum of ₹20,000,000 after retirement.

The calculation is based on an interest rate of 6.5% and period of 25 years. FVIFA will be 62.7154 using the formula ((1+r)-n-1/r)(1+r).

Best option

Bhatia can save ₹318,901 at the beginning of every year for the next 25 years and will receive an amount equivalent to ₹20,000,000 at end of year 55.

N/B: Interest rate is assumed to be 6.5% and expected amount is projected at ₹20,000,000